elination method to solve this linear euation: 3x+7y=1
2x+9y=5

Answer by oberobic(2304)

3x +7y = 1
2x + 9y = 5
.
2(3x +7y) = 2(1)
6x + 14y = 2
.
2x +9y = 5
3(2x +9y) = 3(5)
6x +27y = 15
.
Now we're ready to use the elimination method for the 'x' term.
.
6x + 14y = 2
6x +27y = 15
--------------
0x -13y = -13
.
y = 1
.
Now substitute y=1 into one of the equations to find 'x'.
.
3x +7y = 1
3x +7(1) = 1
3x + 7 = 1
3x = -6
.
x = -2
.
Always check your work.
Substitute these values into the other equation to be sure.
.
2x + 9y = 5
2x =2(-2) = -4
9y = 9(1) = 9
-4 + 9 = 5
Correct.
.
Answer: x=-2, y=1
.
Done.